5. A 6-in.-diameter well that fully penetrates a 40-ft-thick open aquifer is pumped at a rate of 100 gpm. The drawdown in the6TH question
Answer #1

Calculation...

- Asper given data : n=0.5 feet Ground level y Water table 14 draw down curve. 3= 22 ft H= 40 feet h4042 - 28 feet Th=40-22 =volution of che Validity of above result... - medium should be homogenous means propertys of soil - in any direction of ever